How could you use the product 10 times 414 = 4,140 to find the product of 12 times 414?
Answer by stanbon(75887)

How could you use the product 10 times 414 = 4,140 to find the product of 12 times 414?
---
12(414) = (10+2)(414) = 10*414+2(414) = 4140+828 = 4968
===============
Cheers,
Stan H.